# Learning to Generate Novel Scientific Directions with Contextualized Literature-based Discovery

### 22 Jan 2024 v4

@@ Appendix E Automated evaluation for idea-node prediction
${}^{\dagger}$ $*{}^{*}$
CSRoBERTa+CT
0.080
0.043
0.111
0.151
T5+Div+CT
0.045
0.032
0.065
0.076
Table 11: Canonicalized KG metrics for idea node prediction on the whole dataset.
E.1 Evaluation Metrics
Evaluating new idea predictions given an existing scientific KG is non-trivial for non-canonicalized KGs such as ours. Traditional metrics such as MRR and Hits, which can be used for the link prediction task in canonicalized KGs, cannot deal with fine-grained/complex ground truth due to IE errors. For example, because coreference resolution cannot correctly capture all coreference relations, models might predict a correct mention that is marked wrong because its coreference with the target entity is undetected. Moreover, given the vastness of the search space for new ideas, numerous plausible ideas may fail to align with the ground truth precisely. Therefore, we introduce two new similarity-based evaluation metrics for natural language generation, which will compute the overlap between ground truth and predicted entities as “soft” Hits. Given the top-10 predictions $p_{i}$, the corresponding reference $r$, and one of the similarity-based metrics such as BERTScore Zhang* et al. (2020), and BARTScore Yuan et al. (2021), we can calculate two versions of the metric:
$\mathrm{AvgM}$ $=\frac{\sum_{i=1}^{10}\frac{\mathrm{metric}(p_{i},r)}{i}}{\sum_{i=1}^{10}\frac{1}{i}}$ (2)
$\mathrm{MaxM}$ $=\max_{i=1,2,...,10}\left(\mathrm{metric}(p_{i},r)\right)$ (3)
Unlike MRR and Hits, these metrics can capture predictions that are semantically related to the ground truth, even if they don’t match on the surface. Because the domain of our link prediction results is computer science, we use SciBERT Beltagy et al. (2019) in BERTScore. Additional details of the evaluation setup are in Appendix C.
The canonicalized evaluation results are shown in Table 11. In Table 13, for each model variant we test, we compare the best dual-encoder with its T5 counterpart. [^19] GPT3.5 is not shown as it consistently underperforms. However, low-quality retrieval neighbors hurt model performance for baselines with relatively high performance, such as dual-encoder models. These observations echo the observations made for the sentence generation task. Different from the previous task, we observe that the in-context contrastive objective has little impact on the dual-encoder models, despite the fact that it shows consistent improvement for T5. The main reason is that the dual-encoder baselines already utilize contrastive learning objectives, while adding additional negatives might not help the model.
Type MRRnormal-$\uparrow$ HIT@1normal-$\uparrow$ HIT@5normal-$\uparrow$ HIT@10normal-$\uparrow$
Forward 0.126 0.073 0.177 0.232
Backward
0.054
0.028
0.072
0.100
Table 12: Canonicalized KG metrics evaluation results on idea node prediction for the whole dataset with SciBERT+CT.
Model Avg BARTnormal-$\uparrow$ Max BARTnormal-$\uparrow$ Avg BERTnormal-$\uparrow$ Max BERTnormal-$\uparrow$
SciBERT
-6.188
-4.762
0.708
0.796
T5+Div
-6.268
-5.065
0.708
0.785
[…]
SciBERT+CL
-6.193
-4.768
0.708
0.796
T5+Div+CL
-6.267
-5.061
0.708
0.785
[…]
SciBERT+SN
-6.209
-4.803
0.708
0.794
T5+Div+SN
-6.261
-5.050
0.709
0.786
[…]
SciBERT+KG
-6.207
-4.823
0.708
0.794
T5+Div+KG
-6.285
-5.079
0.707
0.785
[…]
SciBERT+CT
0.710
T5+Div+CT
-6.191
-4.924
0.793
Table 13: Non-canonicalized KG metrics for idea node prediction on the whole dataset.
Model Avg BARTnormal-$\uparrow$ Max BARTnormal-$\uparrow$ Avg BERTnormal-$\uparrow$ Max BERTnormal-$\uparrow$
SciBERT
-6.783
-5.908
0.675
0.733
T5+Div
-6.750
-5.834
0.678
0.738
[…]
SciBERT+CL
-6.778
-5.913
0.676
0.734
T5+Div+CL
-6.681
-5.817
0.680
0.739
[…]
SciBERT+SN
-6.715
-5.856
0.676
0.733
T5+Div+SN
-6.752
-5.856
0.678
0.737
[…]
SciBERT+KG
-6.764
-5.909
0.675
0.733
T5+Div+KG
-6.750
-5.884
0.676
0.738
[…]
SciBERT+CT
-6.715
-5.820
0.679
0.737
T5+Div+CT -6.680 -5.797 0.682 0.739
Table 14: Non-canonicalized KG metrics evaluation results on idea node prediction for the gold subset.
E.2 Gold evaluation
We annotate a gold dataset similar to §4.2. Because our idea node prediction task heavily relies on IE quality, we add another condition to the existing three criteria: whether the IE is of sufficient quality. The annotation details are in Appendix G. We obtain 107 pairs. The results are in Table 14. Because we only keep instances where the background context and ground truth are not similar, the scores are worse than Table 13. Furthermore, given that our models achieve a low number of correct predictions in this gold subset due to the quality of the information extraction, we decide not to use canonicalized KG metrics. Unfortunately, we find that the results do not show statistical significance between any pairs of outputs.
Model MRRnormal-$\uparrow$ HIT@1normal-$\uparrow$ HIT@3normal-$\uparrow$
SciBERT
0.671
0.434
0.933
T5+CBS
0.522
0.238
0.795
[…]
SciBERT+CL
0.672
0.434
0.934
T5+CBS+CL
0.539
0.261
0.809
[…]
SciBERT+SN
0.670
0.431
0.933
T5+CBS+SN
0.523
0.240
0.793
[…]
SciBERT+KG
0.668
0.429
0.931
T5+CBS+KG
0.518
0.231
0.793
[…]
SciBERT+CT
T5+CBS+CT
0.532
0.256
0.800
Table 15: Multi-choice evaluation results on idea node prediction for the whole dataset.
E.3 Multi-choice evaluation
Furthermore, we conduct a multi-choice evaluation for idea-node prediction to test the models’ ability to distinguish similar candidates. Specifically, for each instance in the test set, we select three nodes that closely resemble the ground truth from the same paper. The selection is based on cosine similarity provided by SentenceBert. We ask the models to distinguish the best candidate out of four candidates, including the ground truth. Because of the constrained output space, we employ T5+CBS for this experiment as the sequence-to-sequence baseline. The sample input and candidates for multi-choice evaluation are in Table 17. The results are in Table 15. The results align closely with Table 11 and 13, suggesting citation neighbors successfully help models generate better results. Moreover, dual encoders are more effective when distinguishing similar candidates.
Moreover, we annotate a subset from the gold subset where all coreference and valid candidates, except the ground truth, are excluded. The results are in Table 16. However, similar to Table 14, the results between different SciBERT models are not statistically significant due to sample size.
Model MRRnormal-$\uparrow$ HIT@1normal-$\uparrow$ HIT@3normal-$\uparrow$
SciBERT
0.756
0.573
0.961
T5+CBS
0.558
0.301
0.845
[…]
SciBERT+CL
0.756
0.573
0.961
T5+CBS+CL
0.538
0.262
0.816
[…]
SciBERT+SN
0.761 0.583
0.961
T5+CBS+SN
0.554
0.301
0.835
[…]
SciBERT+KG
0.748
0.563
0.971
T5+CBS+KG
0.548
0.272
0.816
[…]
SciBERT+CT
0.740
0.553
0.942
T5+CBS+CT
0.568
0.320
0.806
Table 16: Multi-choice evaluation results on idea node prediction for the filtered gold dataset.
Figure 4: Gold subset annotation interface
Type Content
Seed Term Prompt
symbolic reasoning is done by using OtherScientificTerm
[…]
Context
Tables are often created with hierarchies, but existing works on table reasoning mainly focus on flat tables and neglect hierarchical tables. Hierarchical tables challenge numerical reasoning by complex hierarchical indexing, as well as implicit relationships of calculation and semantics.
[…]
Ground Truth
hierarchy-aware logical form
[…]
Distractor
hierarchical structure, hierarchical tables, symbolic reasoning
Table 17: Input and candidates for multi-choice evaluation on idea node prediction
Appendix F Further result analysis
Impact of Retrieved Neighbor Similarity
We also analyze the quality of different types of retrieved neighbors. We first report the number of instances with corresponding neighbors for each dataset in Table 18. We observe that KG provides the least neighbors while SN provides the most neighbors. Because our test set is built from new papers not observed during training, and our non-canonicalized KG is sparse, many nodes might not have any neighbors. Among those instances where all three neighbor types exist, we compute the semantic similarity between retrieved neighbors and the ground truth. The results are in Figure 6. For node prediction, it is evident that the citation neighbors provide the context most relevant to the ground truth, explaining the superior performance of models with citation-based retrieval in Table 11. For sentence generation, we observe that the semantic neighbors are closer to the ground truth, which matches our findings in Table 2. We note that the cosine similarities are not directly comparable across the two settings: the ground truth for node prediction is made of short entity mentions, while the ground truth for sentence generation consists of full sentences. Table 8 shows examples of retrieved neighbors for idea-sentence generation and idea-node prediction. Semantic neighbors tend to be shorter than citation neighbors, making them more similar to the idea-sentence ground truth.
Because the ground truth for sentence generation consists of full sentences, the cosine similarity is naturally lower for citation neighbors than for semantic neighbors since citation neighbors are longer than semantic neighbors (as shown in Table 19); for the same reason, however, they provide more background knowledge and thus still perform the best in Table 9.
Figure 5: Cosine similarity between background context and the target node, for forward/backward edge directions. Backward edges of the form [,used-for,task] benefit less from background context.
Task SN KG CT
Sentence Generation
69
4,122
363
Node Prediction
71
4,306
377
Table 18: Number of instances that do not have any neighbors for the test set.
Figure 6: Cosine similarity between the ground truth and the least similar retrieved neighbor, where all three neighbors exist.
Task SN KG CT
Sentence Generation
3.11
4.09
8.73
Node Prediction
3.12
4.10
8.74
Table 19: Average number of tokens for retrieved neighbors and the target where all three neighbors exist
Type R-Lnormal-$\uparrow$ BARTnormal-$\uparrow$ BERTnormal-$\uparrow$
Forward 0.287 -5.557 0.698
Backward
0.247
-5.681
0.683
Table 20: Automatic evaluation results on idea sentence generation for the whole dataset with T5+CT+CL.
Impact of Predicting Methods vs. Tasks
As a reminder, as part of our construction, we extract salient seed terms from target sentences (§2). The seeds are mentions of scientific concepts that are linked via a “used-for” relation, resulting in tuples of the form [head,used-for,?] (forward direction) or [,used-for,tail] (backward). The head/tail mentions can loosely be considered as methods/tasks; for example, a forward prediction tuple would be [Contrastive Learning Approach,used-for,?]. In the node prediction setting, we aim to predict the missing tail/head. In sentence generation, we generate a full sentence that should include a reference to the missing term.
We explore differences in performance between the two directions. In Table 20, we report the scores for the forward and backward directions, using the best models in the sentence-level settings (see Table 12 for node-level). Forward predictions are better than backward predictions, even though we have more backward training pairs (Table 1). In other words, predicting tasks given methods, is easier than predicting methods given tasks.
One reason for this difference is that background context sentences naturally tend to discuss main problems in specific areas—that is, they discuss tasks that need to be addressed. This type of background information helps inform predictions of the missing tasks in forward-direction tuples [method,used-for,?]. In contrast, backward-direction tuples [?,used-for,task] benefit less from the background, since ideas proposed in papers often suggest new methods to address a given background task. See also Figure 5.
Impact of Retrieval and In-Context Contrastive Objective
We conclude with qualitative observations on the impact of retrieval and our in-context contrastive objective. We observe that the retrieved neighbors enrich the input context with relevant knowledge. For example, given the seed term “multi-modal interactions” as a method, and the cited paper title “cross-modal relationship inference for grounding referring expressions”, SciBERT+CT correctly predicts “referring image segmentation” as the second-highest ranked task (relation extracted from Jain and Gandhi (2022)).
We also confirm qualitatively that the in-context contrastive objective reduces the model’s tendency to copy directly from context. For example, given the seed term “graph neural networks (gnns)” and context “…interest in word alignments is increasing again for their usefulness in domains such as typological research…”, SciBERT directly copies “alignments” as part of the top result, while SciBERT+CL predicts “bilingual graphs” which is closer to the ground truth “multi parallel word alignment”.
@@ Appendix G Gold Dataset Annotation Details
Complex sphingolipids and sterols are coordinately involved in various cellular functions, e.g. the formation of lipid microdomains.
@@
PMC-LLaMA
AbaA domains could be classified into different groups, including GrA, GrB, GrC, GrF, and AbaA domains.
@@
PMC-LLaMA + CT
Among them, the clinical severity is an important question. Furthermore, we would like to explore the effects of AbA on the development of the severity, onset of biogenic factors and the severity/pharmacokinetic drug interaction.
@@
PMC-LLaMA + KG
Furthermore, we found that AbA utilization was significantly reduced in the absence of Bs1.
@@
PMC-LLaMA + SN
When A/N inoculum size) the most abundant PAH on LF is significantly lower than that of Arab1.
Table 23: Input and idea-sentence generation results for seed chemical “AbA” in the biochemical domain.
@@
Appendix H Human Evaluation Details for Idea-Sentence Generation
The instructions for human evaluation can be found in Figure 7, while an example of the human evaluation interface is provided in Figure 8 and  9. Human annotators are required to evaluate each system output based on the following criteria: (1) Is the candidate relevant to the context + seed term? (2) Does the candidate copy too much from the context, or is it sufficiently novel/different from the context? (3) Does the candidate’s suggestion generally make sense to you scientifically? (4) Is the language sufficiently clear and coherent to understand the suggestion? The input for sample human annotation is in Table 21 and the human labels are in Table 22. The human annotation agreement is in Table 24.
@@ Appendix H Human Evaluation Details for Idea-Sentence Generation
Table 24: Percent (%) of same labels from overlapped 10 human evaluation instances on each pair of annotators.
Appendix I Biochemical Case Study
I.1 Dataset Collection
We collect PubMed papers from 1988 to 2023 using Entrez Programming Utilities API202020www.ncbi.nlm.nih.gov/books/NBK25501/. We use PubTator 3 Islamaj et al. (2021); Wei et al. (2022); Luo et al. (2023); Wei et al. (2023); Lai et al. (2023). The PubTator 3 performs named entity recognition, relation extraction, entity coreference and linking, and entity normalization for the abstracts in the dataset. PubTator 3 identifies bio entities belonging to seven types: gene, chemical, chromosome, cell line, variant, disease, and speciesl and relations belonging to 13 types: associate, cause, compare, convert, contract, drug interact, inhibit, interact, negative correlate, positive correlate, prevent, stimulate, and treat. Finally, we use a sentence classifier trained on CODA-19 Huang et al. (2020) to classify sentences in abstracts into background, purpose, method, finding, and other. We select sentences with labels of background as background context and remove sentences with labels of other. We treat the rest sentences that have at least one entity as the target sentence. Similar to the construction of the silver dataset, we only keep samples with low similarity between background context and corresponding ground truth sentences  [^21] The similarity is calculated with all-mpnet-base-v2.. Our final dataset has 6,491 papers before 2023/02, 778 papers from 2023/02 to 2023/04, and 663 papers after 2023/05.
I.2 Training Details
Our PMC-LLaMA and its variants are built based on the Huggingface framework (Wolf et al., 2020)222222github.com/huggingface/transformers. We use its chaoyi-wu/PMC_LLAMA_7B_10_epoch as the base model. We finetune those models with a learning rate of $6\times 10^{-6}$ and $\epsilon=5\times 10^{-7}$. The maximum training epoch for all models is $15$. All models are fine-tuned on 4 NVIDIA A100 80 GB GPUs with Fully Sharded Data Parallel [^23] https://huggingface.co/docs/accelerate/usage_guides/fsdp. The training time for each model is about 20 hours. The input and sample outputs are in Table 23.
Appendix J Scientific Artifacts
We list the licenses of the scientific artifacts used in this paper: Semantic Scholar Academic Graph API (API license agreement242424api.semanticscholar.org/license/), Huggingface Transformers (Apache License 2.0), SBERT (Apache-2.0 license), BARTScore (Apache-2.0 license), BERTScore (MIT license), SimKGC (no license), PMC-LLaMA (no license), Entrez Programming Utilities API (Copyright252525www.ncbi.nlm.nih.gov/books/about/copyright/), PubTator 3 (Data use policy262626www.ncbi.nlm.nih.gov/home/about/policies/), and OpenAI (Terms of use272727openai.com/policies/terms-of-use).
Appendix K Ethical Consideration
The contextualized literature-based discovery task and corresponding models we have designed in this paper are limited to the natural language processing (NLP) domain, and might not apply to other scenarios.
K.1 Usage Requirement
@@ I.2 Training Details
K.2 Data Collection
We collect 67,408 ACL Anthology papers from 1952 to 2022 using Semantic Scholar Academic Graph API, under API license agreement [^28] https://api.semanticscholar.org/license/. We ensure our data collection procedure follows the Terms of Use at https://allenai.org/terms. According to the agreement, our dataset can only be used for non-commercial purposes. As mentioned in §4, we perform the human evaluation. All annotators involved in human evaluation are voluntary participants with a fair wage. We further collect 7,932 PubMed papers from 1988 to 2023 using Entrez Programming Utilities API292929www.ncbi.nlm.nih.gov/books/NBK25501/. We follow their data usage guidelines303030www.ncbi.nlm.nih.gov/books/about/copyright/.

### 17 Feb 2024 v5

@@ B.3 Automated Evaluation
Table 9: Automatic evaluation results for the challenging and gold subsets. CL is a model with in-context contrastive augmentation. SN is a model with semantic inspirations. KG is a model with KG inspirations. CT is a model with citation inspirations. R-L denotes ROUGE-L. BERT denotes BERTscore with SciBERT as its encoder. ${}^{\dagger}$ indicates that differences between models are not statistically significant ($p\leq 0.05$) when compared to each other but are still significant when compared to the other models based on t-test.
@@
Huangjiu is known for its unique aroma, primarily attributed to its high concentration of b-phenylethanol (ranging from 40 to 130 mg/L).
@@
Meditron
We then selected six key genes involved in the phenylalanine metabolism pathway and predicted that their product flux may affect the synthesis of b-phenylethanol.
@@
Meditron + CT
We found that the key amino acid residue that controls the activity of Aro10p was not conserved in wine yeast strains, which may explain the lower b-phenylethanol production in wine fermentation compared with that in Chinese huangjiu.
@@
Meditron + KG
Both target genes, SSA1 and ARO10, were deleted using the CRISPR-Cas9 genome editing system.
@@
Meditron + SN
Herein, we report that the key barrier for b-phenylethanol production in Huangjiu is ARO10, the only bi-functional amino acid decarboxylase in Saccharomyces cerevisiae.
Table 12: Input and idea-sentence generation results for seed gene “ARO10” in the biochemical domain.
@@
Appendix C Human Annotation and Evaluation Details
Gold Dataset Annotation Details
Study I
The instructions for human evaluation can be found in Figure 6, while an example of the human evaluation interface is provided in Figure 7 and  8. Human annotators are required to evaluate each system output based on the following criteria: (1) Is the candidate relevant to the context + seed term? (2) Does the candidate copy too much from the context, or is it sufficiently novel/different from the context? (3) Does the candidate’s suggestion generally make sense to you scientifically? (4) Is the language sufficiently clear and coherent to understand the suggestion? The input for sample human annotation is in Table 10 and the human labels are in Table 11. The human annotation agreement is in Table 13.
Study III
We ask the following questions to human annotators to evaluate the quality of regeneration results: (1) Is the regenerated idea substantially different from the original? (2) Is the regenerated idea more novel and creative than the original idea? (3) Does the second iteration increase novelty? The human annotation agreement is in Table 14.
@@ Appendix C Human Annotation and Evaluation Details
Table 13: Percent (%) of same labels from overlapped 10 human evaluation instances on each pair of annotators for Study I.
Annotator Pair 1-2 1-3 1-4 1-5
Agreement %
92.5
93.3
87.5
90.0
Table 14: Percent (%) of same labels from overlapped 20 human evaluation instances on each pair of annotators for Study III. (1-3) has 60 shared questions. The rest of the pairs each share 40 questions.
Appendix D Scientific Artifacts
We list the licenses of the scientific artifacts used in this paper: Semantic Scholar Academic Graph API (API license agreement222222api.semanticscholar.org/license/), Huggingface Transformers (Apache License 2.0), SBERT (Apache-2.0 license), BERTScore (MIT license), Meditron-7b (Llama2), Entrez Programming Utilities API (Copyright232323www.ncbi.nlm.nih.gov/books/about/copyright/), PubTator 3 (Data use policy242424www.ncbi.nlm.nih.gov/home/about/policies/), and OpenAI (Terms of use252525openai.com/policies/terms-of-use).
Appendix E Ethical Consideration
The SciMON task and corresponding models we have designed in this paper are limited to the natural language processing (NLP) and biochemical domain, and might not apply to other scenarios.
E.1 Usage Requirement
@@
E.2 Data Collection
We collect 67,408 ACL Anthology papers from 1952 to 2022 using Semantic Scholar Academic Graph API, under API license agreement [^26] https://api.semanticscholar.org/license/. We ensure our data collection procedure follows the Terms of Use at https://allenai.org/terms. According to the agreement, our dataset can only be used for non-commercial purposes. As mentioned in §4, we perform the human evaluation. All annotators involved in human evaluation are voluntary participants with a fair wage. We further collect 5,708 PubMed papers from 1988 to 2024 using Entrez Programming Utilities API272727www.ncbi.nlm.nih.gov/books/NBK25501/. We follow their data usage guidelines282828www.ncbi.nlm.nih.gov/books/about/copyright/.